OpenStudy (anonymous):

calculate the derivative of y with respect to x of X^2y+2xy^2=x+y

OpenStudy (anonymous):

it would be the last answer right?

OpenStudy (anonymous):

\[x^2y+2xy^2=x+y\] \[2xy+x^2y'+2y^2+4xyy'=1+y'\] then a raft of algebra to solve for \[y'\]

OpenStudy (anonymous):

if you wanted to use wolfram it would be the very last answer on this page http://www.wolframalpha.com/input/?i=x^2y%2B2xy^2%3Dx%2By

OpenStudy (anonymous):

\[y'(x^2+4xy-1) = (1-2xy-2y^2)\]

OpenStudy (anonymous):

thank you